OpenStudy (anonymous):

write the equation of the tangent line to the curve y=x^3+3x-4 at (2,10). The equation of the tangent line is ?

OpenStudy (danjs):

y = f(x) Equation for a tangent line at a point (a, f(a)) y - f(a) = f '(a)*(x - a)

OpenStudy (danjs):

The derivative, f '(a) is the slope of the tangent line at the point x=a

OpenStudy (danjs):

Calculate the derivative, then plug into that x=2, f ' (2) = ? y - 10 = f '(2)*(x - 2) just find the derivative evaluated at 2 and put it in the equation above for f '(2)

OpenStudy (anonymous):

would the equation be 3x^2-3?

OpenStudy (anonymous):

or would I plug 2 into that equation?

OpenStudy (danjs):

for y=f(x) The derivative is f '(x) = 3x^2 + 3 Evaluate that at x=2, f '(2) will be the slope for your tangent line

OpenStudy (danjs):

If you have a slope, and a point, recall the point slope formula from algebra... point (a , f(a)) slope = f '(a) y - f(a) = f '(a)*(x - a)

OpenStudy (danjs):

I have to run... the derivative f '(x) = 3x^2 + 3 slope of tangent line at point x=2, y=10 is f '(2) = 15 Tangent line y - f(a) = f '(a)*(x - a) y - 10 = 15*(x - 2) rearrange if you want a different form... here is the graph

OpenStudy (danjs):

Ill be back on later probably..

OpenStudy (anonymous):

so the equation becomes y=15x-20 once you simplify?

OpenStudy (stamp):

\[f(x)=x^3+3x-4\]\[f'(x)=3x^2+3\] The slope of the tangent line at point (2,10) is\[f'(2)=15\] \[y=mx+b\]And we know m = 15 for the point (2, 10). Solve for b. y = 15x + b 10 = 15(2) + b b = -20 The equation of the tangent line is\[y=15x-20\]